# Language Models Encode Numbers Using Digit Representations in Base 10

### v1

@@ 1 Introduction
Recently, Zhu et al. (2024) used linear probes to predict the number encoded in a hidden representation, showing high correlation with the expected value. However, the probes exhibited low accuracy, suggesting that a linear representation alone is not sufficient to explain how LLMs can often perform exact numerical operations, such as addition and multiplication. Maltoni and Ferrara (2024) have suggested that LLMs may do arithmetic in “value space”, but then we would expect to see a normally-distributed error pattern, which we will see is not the case in widely-used models. [^1] For additional related work and discussion, see §B.
@@ 2 Model Errors on Numerical Tasks are Scattered Across Digits
We generated 5,000 queries of addition of $N=7$ operands, which sum into a number between 0 and 1000, and calculated the errors of the model on these queries. Figure 2 displays the error distribution, showing that most errors are exact multiples of 10 and 100. Further, when considering the error distributions for any number of operands between 4 and 8, we observe that about 80% of the errors are in a single output digit, which is often not the units digit.
@@ 3.1 Experimental setting
Using the set of probes for some layer $\ell$, we define a function that reconstructs the value of a number $x$ from its representation $\mathbf{h}^{\ell}$. For every digit $i$, define a function ${\texttt{digit}_{i,b}^{\ell}:\mathbb{R}^{d}\to[b]}$ that predicts the value of that digit by applying $\texttt{digit}_{i,b}^{\ell}:=\frac{b}{2\pi}\cdot\text{atan2}(\mathbf{P}_{i,b}^{\ell}\mathbf{h}^{\ell})$.222atan2 computes the two argument arctangent, which we convert from a signed to an unsigned angle between 0 and $2\pi$. Concatenating the outputs of the functions for all the digits of $x$ provides an estimation of its value in base $b$. For example, the value of a 3-digit number would be reconstructed in base $b$ from its $\ell$-layer representation by concatenating $[\texttt{digit}_{3,b}^{\ell},\texttt{digit}_{2,b}^{\ell},\texttt{digit}_{1,b}^{\ell}]$.
@@ 3.2 Probing recovers digit values in base 10 but not the whole number value
The highest accuracy of 0.91 for Llama 3 8B and 0.92 for Mistral 7B is achieved when reconstructing the numbers in base 10. Moreover, for all other bases, accuracy is substantially lower, typically not exceeding 0.2. Specifically, classifying the number directly (base 2000) succeeds in only $<1\%$ of the cases, which further shows that the direct circular representation of the value in the hidden space is not accurate enough for arithmetic, similarly to the linear representation mentioned earlier. Interestingly, base 5 also has relatively high accuracy, though significantly below base 10.
@@ 3.3 Modifying a digit representation modifies the whole number value accordingly
Since the digit representations are circular in base 10, if we flip a number’s hidden representation along the two directions of the probe (Eq. 2), we would expect the modified representation to encode the same number but with one digit flipped, i.e. the digit corresponding to the probe will now take a value of $v+5\ (\text{mod }10)$ where $v$ was the original digit value before the intervention. For example (Figure 3), flipping the tens digit in the representation of $375$ is expected to produce a representation of $325$. For more details see §C.
@@ Results
For the hundreds digit, the exact intended result was achieved 15% of the time, while 47% of the results were ’close’ to the intended number, e.g., the digit was changed but with an error of 1 from the intended outcome. These numbers were respectively 10% and 50% for the tens digit, and 15% and 50% for the units digit. As a baseline, using a linear intervention following Zhu et al. (2024), but with the appropriate change to the normalization such that a specific number is targeted instead of a general direction, the exact result is achieved in less than 1% of the cases.
@@ Appendix A Additional Results
A.1 Probing representations of numbers in word form
@@ A.3 Accuracy of circular probes
Appendix B Related Work
Representation of numbers in LLMs
Mechanistic interpretability of arithmetic tasks
There has been much interest in looking into how LLMs may perform arithmetic tasks. Recent work has largely focused on either performing in depth analysis of the algorithms learned by toy models (Maltoni and Ferrara, 2024; Nanda et al., 2023; Quirke and Barez, 2024; Yehudai et al., 2024) or analyzing information flow in trained open source LLMs (Stolfo et al., 2023; Chen et al., 2024). Our work complements these efforts and provided a basis for future works in this avenue, by analyzing the representations of numbers in modern LLMs.
Failures of LLMs on arithmetic tasks
Appendix C Causal Intervention Details

### v2

@@ 1 Introduction
Recently, Zhu et al. (2024) used linear probes to predict the number encoded in a hidden representation, showing high correlation with the expected value. However, the probes exhibited low accuracy, suggesting that a linear representation alone is not sufficient to explain how LLMs can often perform exact numerical operations, such as addition and multiplication. Maltoni and Ferrara (2024) have suggested that LLMs may do arithmetic in “value space”, but then we would expect to see a normally-distributed error pattern, which we will see is not the case in widely-used models.
@@ 2 Model Errors on Numerical Tasks are Scattered Across Digits
We generated 5,000 queries of addition of $N=7$ operands, which sum into a number between 0 and 1000, and calculated the errors of the model on these queries. Figure 2 displays the error distribution, showing that most errors are exact multiples of 10 and 100. Further, when considering the error distributions for any number of operands between 4 and 8, we observe that about 80% of the errors are in a single output digit, which is often not the units digit. A similar error analysis of GPT-4o (OpenAI et al., 2024) on 20-operand addition tasks showed similar trends (§A.1).
@@ 3.1 Experimental setting
Using the set of probes for some layer $\ell$, we define a function that reconstructs the value of a number $x$ from its representation $\mathbf{h}^{\ell}$. For every digit $i$, define a function ${\texttt{digit}_{i,b}^{\ell}:\mathbb{R}^{d}\to[b]}$ that predicts the value of that digit by applying $\texttt{digit}_{i,b}^{\ell}:=\frac{b}{2\pi}\cdot\text{atan2}(\mathbf{P}_{i,b}^{\ell}\mathbf{h}^{\ell})$.111atan2 computes the two argument arctangent, which we convert from a signed to an unsigned angle between 0 and $2\pi$. Concatenating the outputs of the functions for all the digits of $x$ provides an estimation of its value in base $b$. For example, the value of a 3-digit number would be reconstructed in base $b$ from its $\ell$-layer representation by concatenating $[\texttt{digit}_{3,b}^{\ell},\texttt{digit}_{2,b}^{\ell},\texttt{digit}_{1,b}^{\ell}]$.
@@ 3.2 Probing recovers digit values in base 10 but not the whole number value
The highest accuracy of 0.91 for Llama 3 8B and 0.92 for Mistral 7B is achieved when reconstructing the numbers in base 10. Moreover, for all other bases, accuracy is substantially lower, typically not exceeding 0.2, serving as a natural baseline for the base 10 results. Specifically, classifying the number directly (base 2000) succeeds in only $<1\%$ of the cases, which further shows that the direct circular representation of the value in the hidden space is not accurate enough for arithmetic, similarly to the linear representation mentioned earlier. Interestingly, base 5 also has relatively high accuracy, though significantly below base 10.
@@ 3.3 Modifying a digit representation modifies the whole number value accordingly
Since the digit representations are circular in base 10, if we flip a number’s hidden representation along the two directions of the probe (Eq. 2), we would expect the modified representation to encode the same number but with one digit flipped, i.e. the digit corresponding to the probe will now take a value of $v+5\ (\text{mod }10)$ where $v$ was the original digit value before the intervention. For example (Figure 3), flipping the tens digit in the representation of $375$ is expected to produce a representation of $325$. For more details see §B.
@@ Results
For the hundreds digit, the exact intended result was achieved 15% of the time, while 47% of the results were ’close’ to the intended number, e.g., the digit was changed but with an error of 1 from the intended outcome. These numbers were respectively 10% and 50% for the tens digit, and 15% and 50% for the units digit. As a baseline, using a linear intervention following Zhu et al. (2024), but with the appropriate change to the normalization such that a specific number is targeted instead of a general direction, the exact result is achieved in less than 1% of the cases. A random baseline would be replacing the numeric token with another random numeric token in the range of the intervention, leading to a random baseline accuracy of 0.1%.
@@
4 Related Work
Representation of numbers in LLMs
Mechanistic interpretability of arithmetic tasks
There has been much interest in looking into how LLMs may perform arithmetic tasks. Recent work has largely focused on either performing in depth analysis of the algorithms learned by toy models (Maltoni and Ferrara, 2024; Nanda et al., 2023; Quirke and Barez, 2024; Yehudai et al., 2024) or analyzing information flow in trained open source LLMs (Stolfo et al., 2023; Chen et al., 2024). Most recently, Zhou et al. (2024) demonstrated that LLMs utilize Fourier features for arithmetic operations, with distinct roles for low- and high-frequency components. Our work complements these efforts and provides a basis for future work in this avenue, by analyzing the representations of numbers in modern LLMs.
Failures of LLMs on arithmetic tasks
@@ Appendix A Additional Results
A.1 Error patterns of GPT-4o
We conducted an additional error analysis using GPT-4o (OpenAI et al., 2024) on 15-operand addition tasks. Increasing the number of operands was necessary due to the model’s high accuracy on simpler addition problems. The results were consistent with the trends observed in Figure 2, showing the majority of errors are at multiples of 10, as seen in Figure 4. This indicates that the fragmented error distribution identified in smaller models persists in larger models.
Increasing the number of digits instead of the number of operands leads to errors in multiples of 100 and 1,000 as well, showing that the error distribution stays indicative of a fragmented representation also for other digits.
Figure 4: Error distribution in 15 operand addition for GPT-4o.
@@ A.3 Accuracy of circular probes
A.4 Probing representations of numbers in word form
Appendix B Causal Intervention Details